### v1

Abstract
We investigate the following Robin eigenvalue problem
$\left\{\begin{array}[]{ll}-\Delta u=\mu u&\text{in}\,\,B,\\ \partial_{\texttt{n}}u+\alpha u=0&\text{on}\,\,\partial B\end{array}\right$.
on the unit ball of $\mathbb{R}^{N}$. We obtain the complete spectral structure of this problem. In particular, for $\alpha>0$, we find that the first eigenvalue is $k_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$, where $k_{\nu+l,m}$ is the $m$th positive zero of $kJ_{\nu+l+1}(k)-(\alpha+l)J_{\nu+l}(k)$. Moreover, when $\alpha\in(-1,0)$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ where $\widehat{k}_{\nu,1}$ denotes the unique zero of $\alpha I_{\nu}(k)+kI_{\nu+1}(k)$, and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$. Furthermore, for $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly 0. Our conclusions indicate the ratio $\mu_{2}/\mu_{1}$ may be positive, negative or zero according to the suitable ranges of the parameter $\alpha$.
Keywords: Payne-Schaefer conjecture; Robin eigenvalue problem; Spectrum; Bessel function
AMS Subjection Classification(2020): 34L20; 34L10; 35J05; 35P10; 35P20
1 Introduction
Consider the following heat conduction equation
$w_{t}=\Delta w,$
where $w$ denotes the temperature as a function of position $x\in\Omega$ with $\Omega\subseteq\mathbb{R}^{N}$ and time $t$. The radiation of heat from a homogeneous body $\Omega$ with surface $\partial\Omega$ into an infinite medium of constant temperature zero is characterized at the surface by a boundary condition of the form
$\partial_{\texttt{n}}w+\sigma w=0,$
where $\sigma\in\mathbb{R}$ is a physical constant and n is the unit outer normal vector to $\partial\Omega$. This condition indicates that the rate of change of temperature in the direction of the inner normal is proportional to the jump in temperature from the exterior to the interior of the body. For some given initial condition, by the method of separation of variables, one has the following Robin eigenvalue problem
$\left\{\begin{array}[]{ll}-\Delta u=\mu u&\text{in}\,\,\Omega,\\ (1-\beta)\partial_{\texttt{n}}u+\beta u=0&\text{on}\,\,\partial\Omega,\end{array}\right$. (1.1)
where $\beta\in\mathbb{R}$, which is also called the elastically supported membrane. For $N=1$, it is well known that problem (1.1) has and only has a sequence of simple eigenvalues. Moreover, the eigenfunctions corresponding to $k$th eigenvalue have k1k-1 simple zeros in $\Omega$ [7]. For $N=2,3$, when $\Omega$ is a ball, Courant and Hilbert [8] also gave the method for calculating eigenvalues and eigenfunctions to this problem but without specific calculation conclusion. For general domains or some special domains, a large number of mathematicians have studied this problem, and we will not list them one by one here, we refer to [12] and its references.
When $\beta=1$, problem (1.1) degenerates into the following fixed membrane problem
$\left\{\begin{array}[]{ll}-\Delta u=\lambda u&\text{in}\,\,\Omega,\\ u=0&\text{on}\,\,\partial\Omega.\end{array}\right$. (1.2)
Let $\lambda_{n}$ (counting the multiplicity) with $n\in\mathbb{N}$ be the $n$th eigenvalue of (1.2). The spectral theory of fixed membrane problems on the ball or interval has long been systematically established by mathematicians (see, for example, [8] for $N\leq 3$ and [5] for $N\geq 4$). For $N=2$, Payne, Pólya and Weinberger [17] proved that
$\lambda_{n+1}-\lambda_{n}\leq\frac{2}{n}\sum_{i=1}^{n}\lambda_{i},\,\,\,n\in\mathbb{N}.$
This inequality has been extended into any dimension by Hile and Protter [13] via establishing
$\sum_{i=1}^{n}\frac{\lambda_{i}}{\lambda_{n+1}-\lambda_{i}}\geq\frac{nN}{4},\,\,\,n\in\mathbb{N}.$
Yang [19] obtained the following more sharp inequality
$\sum_{i=1}^{n}\left(\lambda_{n+1}-\lambda_{i}\right)\left(\lambda_{n+1}-\left(1+\frac{4}{N}\right)\lambda_{i}\right)\leq 0,\,\,\,n\in\mathbb{N}.$
In particular, for $N=2$ and $n=1$, one has that
$\frac{\lambda_{2}}{\lambda_{1}}\leq 3.$
Payne, Pólya and Weinberger conjectured that one could replace the value $3$ by the value that $\lambda_{2}/\lambda_{1}$ assumes when $\Omega$ is a disk. That is to say,
$\frac{\lambda_{2}}{\lambda_{1}}\leq\frac{\lambda_{2}}{\lambda_{1}}\Big|_{\text{disk}}=\frac{j_{1,1}^{2}}{j_{0,1}^{2}}\approx 2.5387.$
The PPW conjecture has been included in problem lists of Yau [20, problem 77]. Yau commented that the truth of the conjecture means that one can determine whether a drum is circular or not by knowing its first two tones. The constant $3$ has been improved by several mathematicians. Specially, Brands [4] obtained $2.686$; then de Vries [18] obtained $2.658$; and finally, Chiti [6] obtained $2.586$. For any dimension, this conjecture can be stated that
$\frac{\lambda_{2}}{\lambda_{1}}\leq\frac{\lambda_{2}}{\lambda_{1}}\Big|_{n-\text{ball}}=\frac{j_{n/2,1}^{2}}{j_{n/2-1,1}^{2}}.$
This PPW conjecture was proved by Ashbaugh and Benguria [1, 2, 3].
For $\beta\neq 1$, we write $\alpha=\beta/(1-\beta)\in\mathbb{R}$. Let $\mu_{n}$ (counting the multiplicity) with $n\in\mathbb{N}$ be the $n$th eigenvalue of the following elastically supported membrane
$\left\{\begin{array}[]{ll}-\Delta u=\mu u&\text{in}\,\,\Omega,\\ \partial_{\texttt{n}}u+\alpha u=0&\text{on}\,\,\partial\Omega.\end{array}\right$. (1.3)
In 2001, parallel to the fixed membrane problem, for $N=2$, Payne and Schaefer [16] proposed the following conjecture.
Payne-Schaefer conjecture. The ratio $\mu_{2}/\mu_{1}$ achieves its maximum for the disk for all values of $\alpha$ or for a range of values of $\alpha$.
They proved that $\mu_{2}/\mu_{1}\leq 3$ for $\alpha\geq\alpha_{*}$ with some positive constant $\alpha_{*}$. In 2003, Henrot [11, Open problem 15] restated this conjecture by proposing the following open problem.
Open problem. For what values of $\alpha$ does the ratio $\mu_{2}/\mu_{1}$ achieve its maximum for the disk?
Although the above question is only for a two-dimensional problem, it is evident that there are similar questions for high-dimensional problems, i.e., whether the ratio $\mu_{2}/\mu_{1}$ achieves its maximum for the ball for all values of $\alpha$ or for a range of values of $\alpha$. Different from the fixed membrane problem, we do not even know the exact value of $\mu_{2}/\mu_{1}$ for the ball. In fact, to the best of our knowledge, we even not know the exact value of $\mu_{2}/\mu_{1}$ in the one-dimensional case. So the main aim of this paper is to determine the exact value of $\mu_{2}/\mu_{1}$ on the ball, which may be the basis for solving the above conjecture or open problem.
From now on, we use $B$ to denote the unit ball of $\mathbb{R}^{N}$ with the center at the origin and assume that $\Omega=B$. In fact, our argument also holds for general balls, but for simplicity, we only consider the unit ball. To realize our main aim, we here provide some more general conclusions.
Theorem 1.1. For problem (1.3) with $\Omega=B$ and $N\geq 2$, there exists a unique sequence of positive eigenvalues $\mu_{l,m}$ for $l\in\mathbb{N}\cup\{0\}$ and $m\in\mathbb{N}$ which has the following properties:
(a) when $\alpha>-l$, the exact value of $\mu_{l,m}$ is $k_{\nu+l,m}^{2}$, where $k_{\nu+l,m}$ is the $m$th positive zero of $kJ_{\nu+l+1}(k)-(\alpha+l)J_{\nu+l}(k)$; for $\alpha=-l$, $\mu_{l,1}=0$ and $\mu_{l,m}$ is $k_{\nu+l,m}^{2}$ with $m\geq 2$; when $\alpha<-l$, $\mu_{l,1}=-\widehat{k}_{\nu+l,1}^{2}$ where $\widehat{k}_{\nu+l,1}$ denotes the unique zero of $\alpha I_{\nu+l}(k)+lI_{\nu+l}(k)+kI_{\nu+l+1}(k)$, and $\mu_{l,m}$ is $k_{\nu+l,m}^{2}$ with $m\geq 2$;
(b) when $\alpha>-l$ or $m\geq 2$, the eigen-subspace corresponding to $\mu_{l,m}$ is generated by
$u_{l,m}=r^{-\nu}J_{\nu+l}\left(k_{\nu+l,m}r\right)G_{l}(\xi);$
for $\alpha=-l$, the eigen-subspace corresponding to $\mu_{l,1}=0$ is generated by $r^{-\alpha}G_{-\alpha}(\xi)$; while, for $\alpha<-l$, the eigen-subspace corresponding to $\mu_{l,1}=-\widehat{k}_{\nu+l,1}^{2}<0$ is generated by $r^{-\nu}I_{\nu+l}\left(\widehat{k}_{\nu+l,1}r\right)G_{l}(\xi)$, where $G_{l}(\xi)$ is any eigenfunction corresponding to $\kappa_{l}=l(l+N-2)$ on N1N-1 dimensional sphere $\mathbb{S}^{N-1}$; in particular, for any $m\geq 2$, $\mu_{0,m}$ is simple and the corresponding eigenfunction is radially symmetric and has exactly m1m-1 simple zeros;
(c) for any fixed $l$, $\mu_{l,m}$ is strictly increasing with respect to $m$ and converges to infinity, meanwhile, for each fixed $l\in\mathbb{N}\cup\{0\}$, when $\alpha\geq-l$ and $m\geq 2$, it is also strictly increasing with respect to $l$;
(d) when $\alpha\geq-l$ or $m\geq 2$, one has that
$0\leq k_{\nu+l,1}<j_{\nu+l,1},\,\,j_{\nu+l+1,m-1}<k_{\nu+l,m}<j_{\nu+l,m}\,\,\text{for}\,\,m\geq 2;$
for $\alpha>0$, the first eigenvalue is $k_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$; when $\alpha\in(-1,0)$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$; for $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly 0; moreover, $k_{\nu,1}$ and $k_{\nu+1,1}$ are strictly increasing with respect to $\alpha$, and as $\alpha$ tends to $\infty$, $k_{\nu,1}$ approaches $j_{\nu,1}$ and $k_{\nu+1,1}$ approaches $j_{\nu+1,1}$.
We mainly use the variable separation method (see, for example, [8]) to obtain the existence of eigenvalues and eigenfunctions. We further utilize some properties of Bessel functions to obtain the desired properties of eigenvalues and the corresponding eigenfunctions. In particular, we obtain the exact value of the second eigenvalue. Indeed, from Theorem 1.1 for $\alpha>0$ we know that
$\frac{\mu_{2}}{\mu_{1}}\Big|_{B}=\frac{k_{\nu+1,1}^{2}}{k_{\nu,1}^{2}}.$
Here we provide approximate values (computed using mathematical software) of $k_{\nu+l,1}$ and the corresponding $\mu_{2}/\mu_{1}$ values (see Table 1) for different values of $\alpha$ and $l$ in the 22D and 33D cases. Note that $\nu=N/2-1$.
$l$ $\nu$ $\alpha=1$ $\alpha=2$ $\alpha=3$ $\alpha=4$ $\alpha=5$ $\alpha=100$ $\alpha=1000$
0 0 1.25578 1.59945 1.78866 1.90808 1.98981 2.38090 2.40242
0 $1/2$ 1.57080 2.02876 2.28893 2.45564 2.57043 3.11019 3.13845
11 0 2.40483 2.73462 2.94960 3.09890 3.20752 3.79360 3.82788
11 1/2 2.74371 3.14159 3.40561 3.59088 3.72638 4.44850 4.48892
$\frac{\mu_{2}}{\mu_{1}}|_{2-\dim B}$ 3.66726 2.92316 2.71938 2.63768 2.59846 2.53875 2.53874
$\frac{\mu_{2}}{\mu_{1}}|_{3-\dim B}$ 3.05095 2.39794 2.21373 2.13832 2.10166 2.04575 2.04575
Table 1: Approximate values of $k_{\nu+l,1}$ and $\mu_{2}/\mu_{1}$ for $l=0,1$ and $N=2,3$.
As can be seen from Table 1, the values of $\mu_{2}/\mu_{1}$ are obviously much smaller than the upper bound $3$ obtained by Payne and Schaefer in the general domain when $\alpha$ is large enough. It can be observed that the value of $\alpha$ influences the value of $\mu_{2}/\mu_{1}$, which justifies the necessity of imposing constraints on $\alpha$ to ensure that $\mu_{2}/\mu_{1}\leq 3$ holds in [16]. This also provides a reference range for improving the upper bounds of Payne and Schaefer.
For $\alpha=-1$, it is obvious that $\frac{\mu_{2}}{\mu_{1}}\Big|_{B}=0$ when $N\geq 2$. For $\alpha\in(-1,0)$ we know that
$\frac{\mu_{2}}{\mu_{1}}\Big|_{B}=-\frac{k_{\nu+1,1}^{2}}{\widehat{k}_{\nu,1}^{2}}.$
Here we also provide approximate values (computed using mathematical software) of $k_{\nu+1,1}$, $\widehat{k}_{\nu,1}$ and the corresponding $\mu_{2}/\mu_{1}$ values (see Table 2) for different values of $\alpha$ in the 22D and 33D cases.
$\nu$ $\alpha=-0.1$ $\alpha=-0.3$ $\alpha=-0.5$ $\alpha=-0.7$ $\alpha=-0.9$
$k_{\nu+1,1}$ 0 1.76104 1.57883 1.35660 1.06842 0.62721
$k_{\nu+1,1}$ $1/2$ 1.98891 1.77934 1.52553 1.19873 0.70207
$\widehat{k}_{\nu,1}$ 0 0.45286 0.80454 1.06569 1.29403 1.50599
$\widehat{k}_{\nu,1}$ 1/2 0.55323 0.97767 1.28784 1.55477 1.79867
$\frac{\mu_{2}}{\mu_{1}}|_{2-\dim B}$ 15.12204-15.12204 3.85102-3.85102 1.62047-1.62047 0.68170-0.68170 0.17345-0.17345
$\frac{\mu_{2}}{\mu_{1}}|_{3-\dim B}$ 12.92465-12.92465 3.31233-3.31233 1.40319-1.40319 0.59444-0.59444 0.15236-0.15236
Table 2: Approximate values of $k_{\nu+l,1}$, $\widehat{k}_{\nu,1}$ and $\mu_{2}/\mu_{1}$ for $N=2$ and $3$.
2 Spectrum of Robin problem in one dimension
To further study the spectrum of Robin problem on the ball, we first consider the one-dimensional case. For $\Omega=(0,1)$, problem (1.3) simplifies to
$\left\{\begin{array}[]{ll}-u^{\prime\prime}=\mu u,\,\,x\in(0,1),\\ -u^{\prime}(0)+\alpha u(0)=0,\\ u^{\prime}(1)+\alpha u(1)=0.\end{array}\right$. (2.1)
It is well known [8, 9] that problem (2.1) has and only has a sequence of simple eigenvalues $\mu_{k}$ with $\mu_{1}<\mu_{2}<\cdots\rightarrow+\infty$. For $\mu\geq 0$, Bucur, Freitas and Kennedy [12, Chapter 4] gave the following relation of eigenvalue $\mu$ and the parameter $\alpha$
$\alpha^{2}+2\alpha\sqrt{\mu}\cot\left(\sqrt{\mu}\right)-\mu=0$
or the equivalent form
$\alpha_{\pm}=-\sqrt{\mu}\cot\left(\sqrt{\mu}\right)\pm\sqrt{\mu\csc^{2}\left(\sqrt{\mu}\right)}.$
Meanwhile, for $\mu\leq 0$, they also obtained that
$\alpha_{+}=-\sqrt{-\mu}\tanh\left(\frac{\sqrt{-\mu}}{2}\right)$
and
$\alpha_{-}=-\sqrt{-\mu}\coth\left(\frac{\sqrt{-\mu}}{2}\right).$
By analyzing the asymptotic behavior of $\alpha$ with respect to $\mu$, they obtained a rough image of the eigenvalue $\mu$ with respect to the parameter $\alpha$.
We here further analyze these relationships to obtain the exact values of eigenvalues, the number of negative eigenvalues, and the specific expressions of the eigenfunctions.
Proposition 2.1. For problem (1.3) with $\Omega=B$, there exists a unique sequence of positive eigenvalues $\mu_{m}$ for $m\in\mathbb{N}$ which has the following properties:
(a) $\mu_{1}>0$ when $\alpha>0$, $\mu_{1}=0$ for $\alpha=0$, $\mu_{1}<0$ and $\mu_{2}>0$ when $\alpha\in(-2,0)$, $\mu_{1}<0$ and $\mu_{2}=0$ for $\alpha=-2$, $\mu_{2}<0$ and $\mu_{3}>0$ when $\alpha<-2$. Moreover, one has that $(m-1)^{2}\pi^{2}<\mu_{m}<m^{2}\pi^{2}$ with $m\geq 1$ for $\alpha>0$, $\mu_{m}=(m-1)^{2}\pi^{2}$ with $m\geq 1$ for $\alpha=0$, $(m-2)^{2}\pi^{2}<\mu_{m}<(m-1)^{2}\pi^{2}$ with $m\geq 2$ for $\alpha\in(-2,0)$ and $(m-2)^{2}\pi^{2}<\mu_{m}<(m-1)^{2}\pi^{2}$ with $m\geq 3$ for $\alpha\leq-2$;
(b) when $\alpha\geq 0$, the exact values of $\mu_{m}$ can be solved from
$\alpha=-k\frac{\cos k}{\sin k}+k\frac{1}{|\sin k|},$
while, when $\alpha\leq 0$, the exact values of $\mu_{m}$ can be solved from
$\alpha=-k\frac{\cos k}{\sin k}-k\frac{1}{|\sin k|}$
or
$\alpha=-\sqrt{-\mu}\tanh\left(\frac{\sqrt{-\mu}}{2}\right)\,\,\text{or}\,\,\alpha=-\sqrt{-\mu}\coth\left(\frac{\sqrt{-\mu}}{2}\right);$
(c) for $\alpha>0$ the eigen-subspace corresponding to $\mu_{m}$ is generated by
$u_{m}=\frac{\sqrt{\mu_{m}}}{\alpha}\cos\left(\sqrt{\mu_{m}}x\right)+\sin\left(\sqrt{\mu_{m}}x\right);$
for $\alpha=0$ the eigen-subspace corresponding to $\mu_{m}$ is generated by $\cos((m-1)\pi x)$; for $\alpha\in(-2,0)$ the eigen-subspace corresponding to $\mu_{1}<0$ is generated by
$\widetilde{u}_{1}=e^{-\sqrt{-\mu_{1}}x}-\frac{\alpha+\sqrt{-\mu_{1}}}{\alpha-\sqrt{-\mu_{1}}}e^{\sqrt{-\mu_{1}}x}$
and the eigen-subspace corresponding to $\mu_{m}$ with $m\geq 2$ is generated by $u_{m}$; for $\alpha=-2$ the eigen-subspace corresponding to $\mu_{1}<0$ is generated by the above $\widetilde{u}_{1}$, corresponding to $\mu_{2}=0$ is generated by 12x1-2x and corresponding to $\mu_{m}$ with $m\geq 3$ is generated by $u_{m}$; for $\alpha<-2$ the eigen-subspace corresponding to $\mu_{m}$ with $m=1,2$ is generated by
$\widetilde{u}_{m}=e^{-\sqrt{-\mu_{m}}x}-\frac{\alpha+\sqrt{-\mu_{m}}}{\alpha-\sqrt{-\mu_{m}}}e^{\sqrt{-\mu_{m}}x}$
and corresponding to $\mu_{m}$ with $m\geq 3$ is generated by $u_{m}$.
Proof. For $\mu\geq 0$, let $k=\sqrt{\mu}$. We have known that
$\alpha_{\pm}=-k\frac{\cos k}{\sin k}\pm k\frac{1}{|\sin k|}.$
When $k\in(0,\pi)$, one sees that
$\alpha_{\pm}=-k\frac{\cos k}{\sin k}\pm k\frac{1}{\sin k}.$
Since
$\lim_{k\rightarrow\pi^{-}}\frac{1-\cos k}{\sin k}=+\infty,$
we find that $\lim_{k\rightarrow\pi^{-}}\alpha_{+}=+\infty$. While, since
$\lim_{k\rightarrow 0^{+}}\frac{1-\cos k}{\sin k}=0,$
we find that $\lim_{k\rightarrow 0^{+}}\alpha_{+}=0$. By some calculations we get that
$\alpha_{+}^{\prime}(k)=\frac{(1-\cos k)(k+\sin k)}{\sin^{2}k}$
for $k\in(0,\pi)$, which implies that $\alpha_{+}$ is strictly increasing in $(0,\pi)$ which starts from 0 and approaches to $\infty$ as $k\rightarrow\pi^{-}$.
About $\alpha_{-}$ in $(0,\pi)$, we have that
$\lim_{k\rightarrow\pi^{-}}\alpha_{-}(k)=-\pi\lim_{k\rightarrow\pi^{-}}\frac{1+\cos k}{\sin k}=0$
and
$\lim_{k\rightarrow 0^{+}}\alpha_{-}(k)=-\lim_{k\rightarrow 0^{+}}k\frac{1+\cos k}{\sin k}=-2.$
Moreover, we can calculate that
$\alpha_{-}^{\prime}(k)=\frac{(1+\cos k)(k-\sin k)}{\sin^{2}k}$
for $k\in(0,\pi)$, which implies that $\alpha_{-}$ is strictly increasing in $(0,\pi)$ which starts from 2-2 and joins to 0.
By repeating the above analysis process, for each $m\in\mathbb{N}$, we find that $\alpha_{+}$ is strictly increasing in $(m\pi,(m+1)\pi)$, starts from 0 and approaches $\infty$ as $k\rightarrow((m+1)\pi)^{-}$, and that $\alpha_{-}$ is also strictly increasing in $(m\pi,(m+1)\pi)$, starts from -$\infty$ and joins to 0 as $k\rightarrow((m+1)\pi)^{-}$.
For $\mu\leq 0$, we have known that
$\alpha_{+}=-\sqrt{-\mu}\tanh\left(\frac{\sqrt{-\mu}}{2}\right)$
and
$\alpha_{-}=-\sqrt{-\mu}\coth\left(\frac{\sqrt{-\mu}}{2}\right).$
Differentiating $\alpha_{+}$ and $\alpha_{-}$ with respect to $\mu$ respectively, we have that
$\alpha_{+}^{\prime}=\frac{1}{2\sqrt{-\mu}}\left[\tanh\left(\frac{\sqrt{-\mu}}{2}\right)+\frac{\sqrt{-\mu}}{2}\text{sech}^{2}\left(\frac{\sqrt{-\mu}}{2}\right)\right]>0$
and
$\alpha_{-}^{\prime}$ $=\frac{1}{2\sqrt{-\mu}}\left[\coth\left(\frac{\sqrt{-\mu}}{2}\right)-\frac{\sqrt{-\mu}}{2}\text{csch}^{2}\left(\frac{\sqrt{-\mu}}{2}\right)\right]$
$=\frac{1}{4\sqrt{-\mu}}\left(\sinh\left(\sqrt{-\mu}\right)-\sqrt{-\mu}\right)\text{csch}^{2}\left(\frac{\sqrt{-\mu}}{2}\right)>0$
for $\mu<0$. Thus we find that both $\alpha_{+}$ and $\alpha_{-}$ are strictly monotonically increasing, $\alpha_{+}>\alpha_{-}$ and $\alpha_{+}$ starts from -$\infty$ and joins to 0 while $\alpha_{-}$ starts from -$\infty$ and joins to 2-2 in $(-\infty,0]$.
Based on the properties of $\alpha_{\pm}$ described above, we provide a schematic diagram (see Figure 1 and note that in this figure, $k=\sqrt{\mu}\geq 0$ for $\mu\geq 0$ and $k=-\sqrt{-\mu}\leq 0$ for $\mu\leq 0$).
Figure 1: The schematic diagram of $\alpha_{+}$ (solid lines) and $\alpha_{-}$ (dotted lines).
We see that $\mu_{1}>0$ when $\alpha>0$, $\mu_{1}=0$ for $\alpha=0$, $\mu_{1}<0$ and $\mu_{2}>0$ when $\alpha\in(-2,0)$, $\mu_{1}<0$ and $\mu_{2}=0$ for $\alpha=-2$, $\mu_{2}<0$ and $\mu_{3}>0$ when $\alpha<-2$, the range of positive $\mu_{m}$ is obvious, which is the Property (a). Further, Property (b) can be also derived from the above properties of $\alpha_{\pm}$. Property (c) can be obtained by the classical Euler undetermined function method.∎
From the monotonicity of $\alpha_{\pm}$ and its asymptotic behavior we can derive that $\mu_{1}$ and $\mu_{2}$ tend to -$\infty$ as $\alpha\rightarrow$-$\infty$. In particular, for $\alpha=-3$, from Proposition 2.1 we obtain that $\mu_{1}\approx-10.52118$ and $\mu_{2}\approx-6.63412$.
3 Spectrum of Robin problem on the ball
In this section, we establish the spectral structure of the Robin eigenvalue problem on the ball. For $N=2$ and $\mu\geq 0$, Bucur, Freitas and Kennedy [12, Chapter 4] derived the following transcendental relation between $\alpha$ and $\mu$
$\alpha J_{k}\left(\sqrt{\mu}\right)+\frac{1}{2}\sqrt{\mu}\left[J_{k-1}\left(\sqrt{\mu}\right)-J_{k+1}\left(\sqrt{\mu}\right)\right]=0$
where $J_{k}$ is the Bessel function. For $\mu<0$, they also obtained that
$\alpha=-\frac{\sqrt{-\mu}\left[I_{k-1}\left(\sqrt{-\mu}\right)+I_{k+1}\left(\sqrt{-\mu}\right)\right]}{2I_{k}\left(\sqrt{-\mu}\right)},$
where $I_{k}$ is the modified Bessel function. Through the above relations, they further obtained the asymptotic expansion of $\mu$ or $\alpha$.
On the basis of the above classic conclusions, we further provide the exact values of eigenvalues on the ball with any $N\geq 2$, as well as several properties of their corresponding eigenfunctions.
Proof of Theorem 1.1. Let $u$ be the eigenfunction corresponding to eigenvalue $\mu$. If $u$ has the form $v(r)G(\xi)$ for $r\in[0,1]$ and $\xi\in\mathbb{S}^{N-1}$, from [5] we have that
$r^{1-N}\left(r^{N-1}v^{\prime}\right)^{\prime}G+r^{-2}v\Delta G+\mu vG=0$
for $r\in(0,1)$. It is well known that the distinct eigenvalues of $\mathbb{S}^{N-1}$ are given by $\kappa_{l}=l(l+N-2)$ for each $l\in\mathbb{N}\cup\{0\}$, and the corresponding eigenfunction is $l$-order spherical harmonic function $G_{l}(\xi)$. That is to say, we know that
$\Delta G_{l}+\kappa_{l}G_{l}=0$
and
$v^{\prime\prime}+\frac{N-1}{r}v^{\prime}+\left(\mu-\frac{\kappa_{l}}{r^{2}}\right)v=0.$
For $\mu>0$, define $k>0$ by $\mu=k^{2}$. Let $\tau=kr$ and $v(r)=z(\tau)$, the above equation can be transformed into
$z^{\prime\prime}+\frac{N-1}{\tau}z^{\prime}+\left(1-\frac{l(l+N-2)}{\tau^{2}}\right)z=0.$
Let $J(\tau)=\tau^{\nu}z$. Then $J(\tau)$ satisfies the equation
$J^{\prime\prime}+\frac{1}{\tau}J^{\prime}+\left(1-\frac{(\nu+l)^{2}}{\tau^{2}}\right)J=0,$
the solution of which is $J(\tau)=J_{\nu+l}(\tau)$. It follows that
$v(r)=r^{-\nu}J_{\nu+l}(kr),$
up to a constant factor. From the boundary condition we derive that $v^{\prime}(1)+\alpha v(1)=0$, i.e.,
$\alpha J_{\nu+l}(k)+lJ_{\nu+l}(k)-kJ_{\nu+l+1}(k)=0$
where we use the formula $J_{\nu}^{\prime}(x)=-J_{\nu+1}(x)+{\nu}/{x}J_{\nu}(x)$ (see [15]).
Since $J_{\nu+l}(k)$ and $J_{\nu+l+1}(k)$ cannot vanish simultaneously, we can see that
$\alpha+l-k\frac{J_{\nu+l+1}}{J_{\nu+l}}(k)=0.$
It follows that
$\alpha=k\frac{J_{\nu+l+1}}{J_{\nu+l}}(k)-l:=\widetilde{h}_{\nu+l}(k).$
From the conclusions of [14, Theorem B ] or [10] we deduce that $\widetilde{h}_{\nu+l}(0)=-l$ and $\widetilde{h}_{\nu+l}(k)$ is strictly increasing in $\left(0,j_{\nu+l,1}\right)$, it has poles $j_{\nu+l,m}$ and is also strictly increasing between any two adjacent poles.
We now consider the negative eigenvalue. For $\mu<0$, we let $t=-\mu$. By using separation of variables again we can obtain that
$v(r)=r^{-\nu}I_{\nu+l}\left(\sqrt{t}r\right),$
up to a constant factor. Using the boundary condition we have that
$\alpha I_{\nu+l}(k)+kI_{\nu+l}^{\prime}(k)-\nu I_{\nu+l}(k)=0,$
where $k=\sqrt{t}$. By the formula [15]
$I_{\nu}^{\prime}(x)=I_{\nu+1}(x)+\frac{\nu}{x}I_{\nu}(x),$
we get that
$\alpha I_{\nu+l}(k)+lI_{\nu+l}(k)+kI_{\nu+l+1}(k)=0.$
It follows that
$\alpha=-k\frac{I_{\nu+l+1}}{I_{\nu+l}}(k)-l:=\widehat{h}_{\nu+l}(k).$
From the conclusions of [14, Theorem C ] we deduce that $\widehat{h}_{\nu+l}(0)=-l$ and $\widehat{h}_{\nu+l}(k)$ is strictly decreasing in $\left(0,+\infty\right)$.
Finally, for $\mu=0$, reasoning as the above, we derive that
$v^{\prime\prime}+\frac{N-1}{r}v^{\prime}-\frac{\kappa_{l}}{r^{2}}v=0.$
By using the Euler undetermined function method, we can obtain that
$v(r)=r^{l},$
up to a constant factor. In view of the boundary condition, one has that $l=-\alpha$.
In conclusion, we define
${h}_{\nu+l}(k)=\left\{\begin{array}[]{ll}\widetilde{h}_{\nu+l}(k)&\text{for}\,\,\mu>0,\\ -l&\text{for}\,\,\mu=0,\\ \widehat{h}_{\nu+l}(k)&\text{for}\,\,\mu<0.\end{array}\right$.
Hence, from the above properties of ${h}_{\nu+l}(k)$, we derive that, for each fixed $l\in\mathbb{N}\cup\{0\}$, ${h}_{\nu+l}(k)=\alpha$ has and only has a sequence of roots $k_{\nu+l,m}$ for any $m\in\mathbb{N}$ (which depends on the position of $\alpha$, see following Figure 2).
(a) $l=0$ and $\mu>0$
(b) $l=0$ and $\mu<0$
(c) $l\neq 0$ and $\mu>0$
(d) $l\neq 0$ and $\mu<0$
Figure 2: The schematic diagram of $h_{\nu+l}(k)$.
For $\alpha>-l$, $\mu_{l,m}=k_{\nu+l,m}^{2}$ are eigenvalues of problem (1.3) where $k_{\nu+l,m}=\widetilde{k}_{\nu+l,m}$ denotes the $m$th zero of $\alpha=\widetilde{h}_{\nu+l}(k)$. For $\alpha=-l$, $\mu_{l,1}=0$ and $\mu_{l,m}=k_{\nu+l,m}^{2}>0$ with $m\geq 2$. For $\alpha<-l$, $\mu_{l,1}=-\widehat{k}_{\nu+l,1}^{2}<0$ and $\mu_{l,m}=k_{\nu+l,m}^{2}>0$ with $m\geq 2$ where $k_{\nu+l,1}=\widehat{k}_{\nu+l,1}$ denotes the unique zero of $\alpha=\widehat{h}_{\nu+l}(k)$. So we have proved the Property (a).
When $\alpha>-l$ or $m\geq 2$, using $\mu_{l,m}=k_{\nu+l,m}^{2}$ we get that
$v(r)=r^{-\nu}J_{\nu+l}\left(k_{\nu+l,m}r\right):=v_{l,m}.$
Then
$u_{l,m}:=v_{l,m}(r)G_{l}(\xi)$
is the eigenfunction corresponding to $\mu_{l,m}$. For $\alpha=-l$, the eigen-subspace corresponding to $\mu_{l,1}=0$ is generated by $r^{-\alpha}G_{-\alpha}(\xi)$. While, for $\alpha<-l$, the eigen-subspace corresponding to $\mu_{l,1}=-\widehat{k}_{\nu+l,1}^{2}<0$ is generated by $r^{-\nu}I_{\nu+l}\left(\widehat{k}_{\nu+l,1}r\right)G_{l}(\xi)$, where $G_{l}(\xi)$ is any eigenfunction corresponding to $\kappa_{l}=l(l+N-2)$ on N1N-1 dimensional sphere $\mathbb{S}^{N-1}$.
Similar to that of [5] with obvious changes we know that the function-space, $L$, consisting of the span, in $L^{2}(B)$, of all eigenfunctions of $B$ obtained by the above procedure, is dense in $L^{2}(B)$. So problem (1.3) only has this eigenvalue sequence $\mu_{l,m}$.
In particular, for any $m\geq 2$, the eigen-subspace corresponding to $\mu_{0,m}$ is one-dimensional which is generated by $v_{0,m}(r)$. So $\mu_{0,m}$ is simple and the corresponding eigenfunction $v_{0,m}(r)$ is radially symmetric. It has been known (see, for example, [5] or [7, Theorem 2.1 of Chap. 8]) that $v_{0,m}(r)$ has exactly m1m-1 simple zeros. So, we have proved property (b).
For any fixed $l\in\mathbb{N}\cup\{0\}$ and $\alpha\geq-l$, from the properties of $h_{\nu+l}(k)$, we derive that $k_{\nu+l,m}$ is strictly increasing with respect to $m$, which implies $\mu_{l,m}$ is strictly increasing with respect to $m$. Moreover, since $0\leq k_{\nu+l,1}<j_{\nu+l,1}$, $j_{\nu+l,m}<k_{\nu+l,m+1}<j_{\nu+l,m+1}$ and $j_{\nu+l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$, we obtain that $k_{\nu+l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$ which indicates that $\mu_{l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$. Moreover, from [14, Theorem A] we know that $\widetilde{h}_{\nu+l}(k)$ is strictly decreasing with respect to $l$. Combining this with the monotonicity of $h_{\nu+l}(k)$ with respect to $k$ implies that $k_{\nu+l,m}$ with $\alpha\geq-l$ or $m\geq 2$ is strictly increasing with respect to $l$. Thus, when $\alpha\geq-l$ or $m\geq 2$, one sees that $\mu_{l,m}$ is strictly increasing with respect to $l$. So, property (c) is also verified.
We finally prove property (d). When $\alpha\geq-l$ or $m\geq 2$, from the properties of $h_{\nu+l}$ we can see that
$0\leq k_{\nu+l,1}<j_{\nu+l,1},\,\,j_{\nu+l,m-1}<k_{\nu+l,m}<j_{\nu+l,m}\,\,\text{for}\,\,m\geq 2.$
Furthermore, when $\alpha>0$, since $h_{\nu+l}\left(j_{\nu+l+1,m-1}\right)=-l$, we derive that
$j_{\nu+l+1,m-1}<k_{\nu+l,m}<j_{\nu+l,m}\,\,\text{for}\,\,m\geq 2.$
When $\alpha>0$, from the above properties one sees that
$k_{\nu+1,1}<j_{\nu+1,1}<k_{\nu,2}.$
Therefore, for $\alpha>0$, we obtain that
$\mu_{1}=k_{\nu,1}^{2}>0$
and
$\mu_{2}=k_{\nu+1,1}^{2}.$
Moreover, it is clear from the monotonicity of $h_{\nu+l}$ that $k_{\nu,1}$ and $k_{\nu+1,1}$ are strictly increasing with respect to $\alpha$. As $\alpha$ tends to $\infty$, $k_{\nu,1}$ approaches $j_{\nu,1}$ and $k_{\nu+1,1}$ approaches $j_{\nu+1,1}$.
From the properties of ${h}_{\nu+l}(k)$ we can derive the following conclusions for $\alpha<0$. When $\alpha\in(-1,0)$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ where $\widehat{k}_{\nu,1}$ denotes the unique zero of $\alpha I_{\nu}(k)+kI_{\nu+1}(k)$, and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$ where $k_{\nu+1,1}$ is the first positive zero of $kJ_{\nu+2}(k)-(\alpha+1)J_{\nu+1}(k)$. For $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ where $\widehat{k}_{\nu,1}$ denotes the unique zero of $I_{\nu}(k)+kI_{\nu+1}(k)$, and the second eigenvalue is exactly 0. For any given $n\in\mathbb{N}$ with $n\geq 2$, when $\alpha\in(-n,1-n)$, the first eigenvalue is $\min_{i\in\{0,\ldots,n-1\}}\left\{-\widehat{k}_{\nu+i,1}^{2}\right\}$ where $\widehat{k}_{\nu+i,1}$ denotes the unique zero of $\alpha I_{\nu+i}(k)+iI_{\nu+i}(k)+kI_{\nu+i+1}(k)$. For $n\in\mathbb{N}$ with $n\geq 2$ and $\alpha=-n$, the first eigenvalue is $\min_{i\in\{0,\ldots,n-1\}}\left\{-\widehat{k}_{\nu+i,1}^{2}\right\}$ where $\widehat{k}_{\nu+i,1}$ denotes the unique zero of $nI_{\nu+i}(k)+iI_{\nu+i}(k)+kI_{\nu+i+1}(k)$. So we obtain the desired conclusions of (d). ∎
From the argument of Theorem 1.1, we also obtain $n$ negative eigenvalues $\widehat{k}_{\nu+i,1}^{2},i\in\{0,\ldots,n-1\}$ when $\alpha\in(-n,1-n)$ with $2\leq n\in\mathbb{N}$. While, for $\alpha=-n$ with $2\leq n\in\mathbb{N}$, besides $n$ negative eigenvalues, 0 is also an eigenvalue. Undoubtedly, the smallest among these $n$ negative eigenvalues is the first eigenvalue, followed by the second eigenvalue. But since we cannot accurately sort them by size, we cannot determine who is the first two, which is fundamentally different from one-dimensional problem (see Proposition 2.1). We conjecture that $\widehat{k}_{\nu+i,1}^{2}$ is decreasing with respect to $i$. If this conjecture was valid, the first eigenvalue is just $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is $\widehat{k}_{\nu+1,1}^{2}$. Note that this conjecture can be derived from the monotonicity of $\widehat{h}_{\nu+l}(k)$ with respect to $l$. Thus, we also conjecture that $\widehat{h}_{\nu+l}(k)$ is decreasing with respect to $l$.
Since $I_{\nu+1}(k)/I_{\nu}(k)$ increases to $1$ as $k$ increases to $\infty$, we find that $\widehat{h}_{\nu+l}(k)\rightarrow$-$\infty$ (monotonously) as $k\rightarrow+\infty$. This implies that $\mu_{l,1}\rightarrow$-$\infty$ as $\alpha\rightarrow$-$\infty$. In particular, the first eigenvalue tends (monotonously) to -$\infty$ as $\alpha\rightarrow$-$\infty$.
Anyway, we have obtained the complete spectral structure of problem (1.3) by utilizing the properties of Bessel functions. In particular, we obtain the exact values of eigenvalues and the exact expressions for the basis of the eigen-subspace. These conclusions themselves are also interesting.
The conflicts of interest statement and Data Availability statement.
There is not any conflict of interest. Data sharing not applicable to this article as no datasets were generated or analysed during the current study.

### v2

@@
 We obtain the complete spectral structure of this problem.
-In particular, for $\alpha>0$, we find that the first eigenvalue is $k_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$, where $k_{\nu+l,m}$ is the $m$th positive zero of $kJ_{\nu+l+1}(k)-(\alpha+l)J_{\nu+l}(k)$.
-Moreover, when $\alpha\in(-1,0)$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ where $\widehat{k}_{\nu,1}$ denotes the unique zero of $\alpha I_{\nu}(k)+kI_{\nu+1}(k)$, and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$.
-Furthermore, for $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly 0.
-Our conclusions indicate the ratio $\mu_{2}/\mu_{1}$ may be positive, negative or zero according to the suitable ranges of the parameter $\alpha$.
+In particular, for $\alpha>0$, the first eigenvalue is $k_{\nu,1}^{2}$ and the second eigenvalue is $k_{\nu+1,1}^{2}$, where $k_{\nu+l,m}$ is the $m$th positive zero of $kJ_{\nu+l+1}(k)-(\alpha+l)J_{\nu+l}(k)$.
+Moreover, when $\alpha\in(-l,1-l)$ with any $l\in\mathbb{N}$, one has $l$ negative (strictly increasing) eigenvalues $\widehat{k}_{\nu+i,1}^{2}$ with $i\in\{0,\ldots,l-1\}$ where $\widehat{k}_{\nu+l,1}$ denotes the unique zero of $\alpha I_{\nu+l}(k)+lI_{\nu+l}(k)+kI_{\nu+l+1}(k)$; while, for $\alpha=-l$, besides $l$ negative (increasing) eigenvalues, 0 is also an eigenvalue.
 Keywords: Payne-Schaefer conjecture; Robin eigenvalue problem; Spectrum; Bessel function
@@
 $\partial_{\texttt{n}}w+\sigma w=0,$
-where $\sigma\in\mathbb{R}$ is a physical constant and n is the unit outer normal vector to $\partial\Omega$.
+where $\sigma$ is a positive physical constant and n is the unit outer normal vector to $\partial\Omega$.
 This condition indicates that the rate of change of temperature in the direction of the inner normal is proportional to the jump in temperature from the exterior to the interior of the body.
@@
 where $\beta\in\mathbb{R}$, which is also called the elastically supported membrane.
+This problem with positive boundary parameter is also related to the wave equation to elastically restoring boundaries.
+Negative Robin parameter is arising in a model for surface superconductivity [15].
 For $N=1$, it is well known that problem (1.1) has and only has a sequence of simple eigenvalues.
@@
 $\frac{\lambda_{2}}{\lambda_{1}}\leq\frac{\lambda_{2}}{\lambda_{1}}\Big|_{\text{disk}}=\frac{j_{1,1}^{2}}{j_{0,1}^{2}}\approx 2.5387.$
-The PPW conjecture has been included in problem lists of Yau [20, problem 77].
+The PPW conjecture has been included in problem lists of Yau [29, problem 77].
 Yau commented that the truth of the conjecture means that one can determine whether a drum is circular or not by knowing its first two tones.
@@
 They proved that $\mu_{2}/\mu_{1}\leq 3$ for $\alpha\geq\alpha_{*}$ with some positive constant $\alpha_{*}$.
-In 2003, Henrot [11, Open problem 15] restated this conjecture by proposing the following open problem.
+In 2003, Henrot [18, Open problem 15] restated this conjecture by proposing the following open problem.
 Open problem. For what values of $\alpha$ does the ratio $\mu_{2}/\mu_{1}$ achieve its maximum for the disk?
 Although the above question is only for a two-dimensional problem, it is evident that there are similar questions for high-dimensional problems, i.e., whether the ratio $\mu_{2}/\mu_{1}$ achieves its maximum for the ball for all values of $\alpha$ or for a range of values of $\alpha$.
-Different from the fixed membrane problem, we do not even know the exact value of $\mu_{2}/\mu_{1}$ for the ball.
-In fact, to the best of our knowledge, we even not know the exact value of $\mu_{2}/\mu_{1}$ in the one-dimensional case.
-So the main aim of this paper is to determine the exact value of $\mu_{2}/\mu_{1}$ on the ball, which may be the basis for solving the above conjecture or open problem.
+Recently, Freitas and Laugesen [13] investigated the first and second eigenvalues for the ball (also see [14] for $2$-dimensional case).
+Then Langford and Laugesen [21] treated the first and second Robin eigenvalues for geodesic disks in Euclidean, hyperbolic and spherical spaces, which does not use special functions.
+When $\alpha>0$, Freitas [10] obtained sharp estimates for the first eigenvalue of the ball and related quantities.
+While, for negative boundary parameter, Antunes, Freitas and Krejčiřík [1] and Freitas and Krejčiřík [12] obtained some bounds for $\mu_{1}$.
+In [22], Laugesen investigated various optimization problems of the first two eigenvalues on rectangles.
+Different from the fixed membrane problem, the value of $\mu_{2}/\mu_{1}$ for the ball is dependent on parameter $\alpha$.
+Here, based on the above important conclusions, we plan to obtain the complete spectrum of the Robin eigenvalue problem on the ball.
+In particular, determine the exact value of $\mu_{2}/\mu_{1}$ on the ball, which may be the basis for solving the above conjecture or open problem.
 From now on, we use $B$ to denote the unit ball of $\mathbb{R}^{N}$ with the center at the origin and assume that $\Omega=B$.
@@
 $u_{l,m}=r^{-\nu}J_{\nu+l}\left(k_{\nu+l,m}r\right)G_{l}(\xi);$
-for $\alpha=-l$, the eigen-subspace corresponding to $\mu_{l,1}=0$ is generated by $r^{-\alpha}G_{-\alpha}(\xi)$; while, for $\alpha<-l$, the eigen-subspace corresponding to $\mu_{l,1}=-\widehat{k}_{\nu+l,1}^{2}<0$ is generated by $r^{-\nu}I_{\nu+l}\left(\widehat{k}_{\nu+l,1}r\right)G_{l}(\xi)$, where $G_{l}(\xi)$ is any eigenfunction corresponding to $\kappa_{l}=l(l+N-2)$ on N1N-1 dimensional sphere $\mathbb{S}^{N-1}$; in particular, for any $m\geq 2$, $\mu_{0,m}$ is simple and the corresponding eigenfunction is radially symmetric and has exactly m1m-1 simple zeros;
+for $\alpha=-l$, the eigen-subspace corresponding to $\mu_{l,1}=0$ is generated by $r^{-\alpha}G_{-\alpha}(\xi)$; while, for $\alpha<-l$, the eigen-subspace corresponding to $\mu_{l,1}=-\widehat{k}_{\nu+l,1}^{2}<0$ is generated by $r^{-\nu}I_{\nu+l}\left(\widehat{k}_{\nu+l,1}r\right)G_{l}(\xi)$, where $G_{l}(\xi)$ is any eigenfunction corresponding to $\kappa_{l}=l(l+N-2)$ on N1N-1 dimensional sphere $\mathbb{S}^{N-1}$; in particular, $\mu_{0,1}$ is simple with positive radial eigenfunction, for any $m\geq 2$, $\mu_{0,m}$ is simple and the corresponding eigenfunction is radially symmetric and has exactly m1m-1 simple zeros;
 (c) for any fixed $l$, $\mu_{l,m}$ is strictly increasing with respect to $m$ and converges to infinity, meanwhile, for each fixed $l\in\mathbb{N}\cup\{0\}$, when $\alpha\geq-l$ and $m\geq 2$, it is also strictly increasing with respect to $l$;
@@
 $0\leq k_{\nu+l,1}<j_{\nu+l,1},\,\,j_{\nu+l+1,m-1}<k_{\nu+l,m}<j_{\nu+l,m}\,\,\text{for}\,\,m\geq 2;$
-for $\alpha>0$, the first eigenvalue is $k_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$; when $\alpha\in(-1,0)$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$; for $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly 0; moreover, $k_{\nu,1}$ and $k_{\nu+1,1}$ are strictly increasing with respect to $\alpha$, and as $\alpha$ tends to $\infty$, $k_{\nu,1}$ approaches $j_{\nu,1}$ and $k_{\nu+1,1}$ approaches $j_{\nu+1,1}$.
-We mainly use the variable separation method (see, for example, [8]) to obtain the existence of eigenvalues and eigenfunctions.
+for $\alpha>0$, the first eigenvalue is $k_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$.
+Moreover, $k_{\nu,1}$ and $k_{\nu+1,1}$ are strictly increasing with respect to $\alpha$, and as $\alpha$ tends to $\infty$, $k_{\nu,1}$ approaches $j_{\nu,1}$ and $k_{\nu+1,1}$ approaches $j_{\nu+1,1}$;
+(e) when $\alpha\in(-l,1-l)$ with any $l\in\mathbb{N}$, one has $l$ negative (strictly increasing) eigenvalues $\widehat{k}_{\nu+i,1}^{2},i\in\{0,\ldots,l-1\}$; while, for $\alpha=-l$ with any $l\in\mathbb{N}$, besides $l$ negative (increasing) eigenvalues, 0 is also an eigenvalue.
+In particular, when $\alpha\in(-1,0)$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$; for $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly 0.
+As mentioned earlier, the existence, and upper and lower bounds estimation of the first two eigenvalues have been obtained in references [1, 13, 12, 10, 14, 21, 22].
+Here the complete spectrum is been obtained, and the precise calculation formulas (through Bessel functions) for the first two are provided.
+Moreover, we can also determine the exact number of negative eigenvalues based on the range of boundary parameter.
+These conclusions can serve as a supplement to the important literature mentioned above and also as a basis for further research on related problems.
+It should be noted that some of the conclusions regarding the first two eigenvalues may have been covered by the aforementioned literatures, but for the sake of completeness, we still list them here.
+We mainly use the variable separation method (see, for example, [9] or [13]) to obtain the existence of eigenvalues and eigenfunctions.
 We further utilize some properties of Bessel functions to obtain the desired properties of eigenvalues and the corresponding eigenfunctions.
-In particular, we obtain the exact value of the second eigenvalue.
+In particular, we obtain the calculation formula for the first two eigenvalues.
 Indeed, from Theorem 1.1 for $\alpha>0$ we know that
@@
 Table 2: Approximate values of $k_{\nu+l,1}$, $\widehat{k}_{\nu,1}$ and $\mu_{2}/\mu_{1}$ for $N=2$ and $3$.
+In [22, Conjecture B], Laugesen conjectured that $\mu_{2}/\mu_{1}$ on bounded Lipschitz domain is decreasing for $\alpha>0$. The data in Tables 1 and 2 supports this conjecture (at least on the ball).
 2 Spectrum of Robin problem in one dimension
@@
 It is well known [9, 16] that problem (2.1) has and only has a sequence of simple eigenvalues $\mu_{k}$ with $\mu_{1}<\mu_{2}<\cdots\rightarrow+\infty$.
-For $\mu\geq 0$, Bucur, Freitas and Kennedy [12, Chapter 4] gave the following relation of eigenvalue $\mu$ and the parameter $\alpha$
+For $\mu\geq 0$, Bucur, Freitas and Kennedy [19, Chapter 4] gave the following relation of eigenvalue $\mu$ and the parameter $\alpha$
 $\alpha^{2}+2\alpha\sqrt{\mu}\cot\left(\sqrt{\mu}\right)-\mu=0$
@@
 Property (c) can be obtained by the classical Euler undetermined function method.∎
+As far as we know, the key value $\alpha=-2$ is our new discovery.
+This point is the key point for the change in the number of negative eigenvalues.
+When $\alpha<-2$, there are exactly two negative eigenvalues, and when $\alpha\in[-2,0)$, there is exactly one negative eigenvalue.
 From the monotonicity of $\alpha_{\pm}$ and its asymptotic behavior we can derive that $\mu_{1}$ and $\mu_{2}$ tend to -$\infty$ as $\alpha\rightarrow$-$\infty$.
 In particular, for $\alpha=-3$, from Proposition 2.1 we obtain that $\mu_{1}\approx-10.52118$ and $\mu_{2}\approx-6.63412$.
+For $\alpha>0$, we have that $\mu_{1}\in(0,\pi)$ and $\mu_{2}\in(\pi,2\pi)$.
+In this case, Freitas and J.B.
+Kennedy [11] obtained more refined upper and lower bound estimates for the first two eigenvalues.
 3 Spectrum of Robin problem on the ball
 In this section, we establish the spectral structure of the Robin eigenvalue problem on the ball.
-For $N=2$ and $\mu\geq 0$, Bucur, Freitas and Kennedy [12, Chapter 4] derived the following transcendental relation between $\alpha$ and $\mu$
-$\alpha J_{k}\left(\sqrt{\mu}\right)+\frac{1}{2}\sqrt{\mu}\left[J_{k-1}\left(\sqrt{\mu}\right)-J_{k+1}\left(\sqrt{\mu}\right)\right]=0$
+For $N=2$ and $\mu\geq 0$, Bucur, Freitas and Kennedy [19, Chapter 4] derived the following transcendental relation between $\alpha$ and $\mu$
+$\alpha J_{k}\left(\sqrt{\mu}\right)+\frac{1}{2}\sqrt{\mu}\left[J_{k-1}\left(\sqrt{\mu}\right)-J_{k+1}\left(\sqrt{\mu}\right)\right]=0,$
 where $J_{k}$ is the Bessel function. For $\mu<0$, they also obtained that
@@
 $\alpha=k\frac{J_{\nu+l+1}}{J_{\nu+l}}(k)-l:=\widetilde{h}_{\nu+l}(k).$
-From the conclusions of [14, Theorem B ] or [10] we deduce that $\widetilde{h}_{\nu+l}(0)=-l$ and $\widetilde{h}_{\nu+l}(k)$ is strictly increasing in $\left(0,j_{\nu+l,1}\right)$, it has poles $j_{\nu+l,m}$ and is also strictly increasing between any two adjacent poles.
+The above equation also derived by Freitas [10, Section 3] or [13, Section 5] for the special case of $l=0$ and $\alpha>0$.
+From the conclusions of [23, Theorem B ] or [17] we deduce that $\widetilde{h}_{\nu+l}(0)=-l$ and $\widetilde{h}_{\nu+l}(k)$ is strictly increasing in $\left(0,j_{\nu+l,1}\right)$, it has poles $j_{\nu+l,m}$ and is also strictly increasing between any two adjacent poles.
 We now consider the negative eigenvalue. For $\mu<0$, we let $t=-\mu$. By using separation of variables again we can obtain that
@@
 $\alpha=-k\frac{I_{\nu+l+1}}{I_{\nu+l}}(k)-l:=\widehat{h}_{\nu+l}(k).$
-From the conclusions of [14, Theorem C ] we deduce that $\widehat{h}_{\nu+l}(0)=-l$ and $\widehat{h}_{\nu+l}(k)$ is strictly decreasing in $\left(0,+\infty\right)$.
+From the conclusions of [23, Theorem C ] we deduce that $\widehat{h}_{\nu+l}(0)=-l$ and $\widehat{h}_{\nu+l}(k)$ is strictly decreasing in $\left(0,+\infty\right)$.
 Finally, for $\mu=0$, reasoning as the above, we derive that
@@
 $v(r)=r^{l},$
-up to a constant factor. In view of the boundary condition, one has that $l=-\alpha$.
+up to a constant factor. In view of the boundary condition, one has that $l=-\alpha$, which also have been obtained in [13, Section 5].
 In conclusion, we define
@@
 So problem (1.3) only has this eigenvalue sequence $\mu_{l,m}$.
-In particular, for any $m\geq 2$, the eigen-subspace corresponding to $\mu_{0,m}$ is one-dimensional which is generated by $v_{0,m}(r)$.
+In particular, $\mu_{0,1}$ is simple with positive radial eigenfunction, and for any $m\geq 2$, the eigen-subspace corresponding to $\mu_{0,m}$ is one-dimensional which is generated by $v_{0,m}(r)$.
 So $\mu_{0,m}$ is simple and the corresponding eigenfunction $v_{0,m}(r)$ is radially symmetric.
-It has been known (see, for example, [5] or [7, Theorem 2.1 of Chap.
+It has been known (see, for example, [6] or [8, Theorem 2.1 of Chap.
 8]) that $v_{0,m}(r)$ has exactly m1m-1 simple zeros.
@@
 Moreover, since $0\leq k_{\nu+l,1}<j_{\nu+l,1}$, $j_{\nu+l,m}<k_{\nu+l,m+1}<j_{\nu+l,m+1}$ and $j_{\nu+l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$, we obtain that $k_{\nu+l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$ which indicates that $\mu_{l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$.
-Moreover, from [14, Theorem A] we know that $\widetilde{h}_{\nu+l}(k)$ is strictly decreasing with respect to $l$.
+Moreover, from [23, Theorem A] we know that $\widetilde{h}_{\nu+l}(k)$ is strictly decreasing with respect to $l$.
 Combining this with the monotonicity of $h_{\nu+l}(k)$ with respect to $k$ implies that $k_{\nu+l,m}$ with $\alpha\geq-l$ or $m\geq 2$ is strictly increasing with respect to $l$.
@@
 As $\alpha$ tends to $\infty$, $k_{\nu,1}$ approaches $j_{\nu,1}$ and $k_{\nu+1,1}$ approaches $j_{\nu+1,1}$.
+So we obtain the desired conclusions of (d).
+Finally we show property (e).
 From the properties of ${h}_{\nu+l}(k)$ we can derive the following conclusions for $\alpha<0$.
@@
 For $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ where $\widehat{k}_{\nu,1}$ denotes the unique zero of $I_{\nu}(k)+kI_{\nu+1}(k)$, and the second eigenvalue is exactly 0.
-For any given $n\in\mathbb{N}$ with $n\geq 2$, when $\alpha\in(-n,1-n)$, the first eigenvalue is $\min_{i\in\{0,\ldots,n-1\}}\left\{-\widehat{k}_{\nu+i,1}^{2}\right\}$ where $\widehat{k}_{\nu+i,1}$ denotes the unique zero of $\alpha I_{\nu+i}(k)+iI_{\nu+i}(k)+kI_{\nu+i+1}(k)$.
-For $n\in\mathbb{N}$ with $n\geq 2$ and $\alpha=-n$, the first eigenvalue is $\min_{i\in\{0,\ldots,n-1\}}\left\{-\widehat{k}_{\nu+i,1}^{2}\right\}$ where $\widehat{k}_{\nu+i,1}$ denotes the unique zero of $nI_{\nu+i}(k)+iI_{\nu+i}(k)+kI_{\nu+i+1}(k)$.
-So we obtain the desired conclusions of (d).
+In general, for any given $l\in\mathbb{N}$ with $l\geq 2$, when $\alpha\in(-l,1-l)$, the first eigenvalue is $\min_{i\in\{0,\ldots,l-1\}}\left\{-\widehat{k}_{\nu+i,1}^{2}\right\}$ where $\widehat{k}_{\nu+i,1}$ denotes the unique zero of $\alpha I_{\nu+i}(k)+iI_{\nu+i}(k)+kI_{\nu+i+1}(k)$.
+While, for $l\in\mathbb{N}$ with $l\geq 2$ and $\alpha=-l$, the first eigenvalue is $\min_{i\in\{0,\ldots,l-1\}}\left\{-\widehat{k}_{\nu+i,1}^{2}\right\}$ where $\widehat{k}_{\nu+i,1}$ denotes the unique zero of $lI_{\nu+i}(k)+iI_{\nu+i}(k)+kI_{\nu+i+1}(k)$.
+From [13, Lemma 10] we derive that $\widehat{h}_{\nu+l}(k)$ is strictly decreasing with respect to $l$.
+This implies that $\widehat{k}_{\nu+i,1}^{2}$ is strictly decreasing with respect to $i$.
+Hence, $\widehat{k}_{\nu+i,1}^{2}$ is strictly increasing with respect to $i$.
+In particular, the first eigenvalue is just $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is $\widehat{k}_{\nu+1,1}^{2}$.
 ∎
-From the argument of Theorem 1.1, we also obtain $n$ negative eigenvalues $\widehat{k}_{\nu+i,1}^{2},i\in\{0,\ldots,n-1\}$ when $\alpha\in(-n,1-n)$ with $2\leq n\in\mathbb{N}$.
-While, for $\alpha=-n$ with $2\leq n\in\mathbb{N}$, besides $n$ negative eigenvalues, 0 is also an eigenvalue.
-Undoubtedly, the smallest among these $n$ negative eigenvalues is the first eigenvalue, followed by the second eigenvalue.
-But since we cannot accurately sort them by size, we cannot determine who is the first two, which is fundamentally different from one-dimensional problem (see Proposition 2.1).
-We conjecture that $\widehat{k}_{\nu+i,1}^{2}$ is decreasing with respect to $i$.
-If this conjecture was valid, the first eigenvalue is just $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is $\widehat{k}_{\nu+1,1}^{2}$.
-Note that this conjecture can be derived from the monotonicity of $\widehat{h}_{\nu+l}(k)$ with respect to $l$.
-Thus, we also conjecture that $\widehat{h}_{\nu+l}(k)$ is decreasing with respect to $l$.
+From Theorem 1.1 we can see that there are infinite negative eigenvalues when $\alpha\rightarrow$-$\infty$, which is fundamentally different from one-dimensional problem (see Proposition 2.1).
 Since $I_{\nu+1}(k)/I_{\nu}(k)$ increases to $1$ as $k$ increases to $\infty$, we find that $\widehat{h}_{\nu+l}(k)\rightarrow$-$\infty$ (monotonously) as $k\rightarrow+\infty$.